Circular No. 7887

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cp IN NGC 3074 M. Ganeshalingam and W. D. Li, University of California at Berkeley,

report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 28.2 (mag ∼ 17.9) and 30.2 UT (mag ∼ 17.7). SN 2002cp is located at α = 9^{h}59^{m}35^{s}.95, δ = +35^{o}23′57′′.4 (equinox 2000.0), which is 64′′.1 west and 23′′.2 north of the nucleus of NGC 3074. KAIT images taken on Apr. 11.2 (limiting mag ∼ 18.0) and Apr. 3.2 (limiting mag ∼ 19.0) showed nothing at this position.

V4741 SAGITTARII E. Gosset, Institut d’Astrophysique, Liège, writes that a high-resolu-

tion spectrum (range 380–900 nm; resolution 48000) of V4741 Sgr, acquired on Apr. 19.4 UT using the European Southern Observatory 1.5-m telescope (+ FEROS) at La Silla, shows the following full widths at zero intensity (FWZI) and at half-maximum (FWHM) for the lower Balmer lines: Hα, FWZI = 4100 km/s, FWHM = 2800 km/s; Hβ, FWZI = 3850 km/s; Hγ, FWZI = 3800 km/s. Hβ and the upper members of the series are comparatively faint, suggesting marked reddening (cf. IAUC 7879). The interstellar absorption lines related to the Na resonance doublet exhibit a rather complex structure that can be partitioned into three families. The first one is composed of a broad, damped line at ∼ −2 km/s, blended with a narrow line at ∼ −30 km/s. The second family is composed of a narrow line at ∼ −107 km/s with a faint line in the blue wing. The third family is made of three partly blended lines at ∼ −143, −150, and −160 km/s. A fourth family at ∼ −190 km/s could possibly be present. The first three families are also visible, with less resolution, in the Ca^{+} H and K doublet. Gosset adds: “The existence of features up to such velocities is interesting. If these lines were considered as being due to interstellar material showing galactic rotation, we could conclude that the object is situated at a distance quite comparable to that of the Galactic center or even beyond, in good agreement with the reported reddening. However, high-velocity shells are known to exist in the Carina region. The first family component at −2 km/s is also detected for K, CH, and CH^{+}.”

Visual magnitude estimates: Apr. 16.606 UT, 9.8 (A. Pearce, Nedlands, W. Australia); 18.134, 10.3 (R. Y. Shida, São Paulo, Brazil); 18.776, 10.6 (Pearce); 19.755, 10.7 (Pearce).

2002 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green